OpenStudy (anonymous):

F(x)=ln(x) g(x)=e^x+1. F(g(x)) must then be ln(e^x+1) right? Can you simplify this?

OpenStudy (dumbcow):

is that e^(x+1) ?

OpenStudy (dumbcow):

e^x and ln(x) are inverse functions so: \[\ln(e^x) = x\]

OpenStudy (anonymous):

yes you are right

OpenStudy (anonymous):

No, g(x) is just e^x+1, as if it was 1+e^x without any parenthesis

OpenStudy (anonymous):

\[f(g(x))= \ln(e ^{x}+1)\]

OpenStudy (anonymous):

let me know if u need furthur help :)

OpenStudy (anonymous):

oh btw you cant simplify further

OpenStudy (anonymous):

or maybe write Near x=0, I suppose you could expand e^x,

OpenStudy (anonymous):

because ln(1) = 0 and also ln(e^x) is also 0 because they are inverse pairs

OpenStudy (anonymous):

So that makes it x+1 in the end. But the reverse, g(f(x)) = e^ln(x)+1 will also become x+1? Right?

OpenStudy (anonymous):

Ah, ok :)

OpenStudy (anonymous):